# VLD: Visual Language Goal Distance for Reinforcement Learning Navigation

### v1

@@ Abstract
Training end-to-end policies from image data to directly predict navigation actions for robotic systems has proven inherently difficult. Existing approaches often suffer from either the sim-to-real gap during policy transfer or a limited amount of training data with action labels. To address this problem, we introduce Vision–Language Distance (VLD) learning, a scalable framework for goal-conditioned navigation that decouples perception learning from policy learning. Instead of relying on raw sensory inputs during policy training, we first train a self-supervised distance-to-goal predictor on internet-scale video data. This predictor generalizes across both image- and text-based goals, providing a distance signal that can be minimized by a reinforcement learning (RL) policy. The RL policy can be trained entirely in simulation using privileged geometric distance signals, with injected noise to mimic the uncertainty of the trained distance predictor. At deployment, the policy consumes VLD predictions, inheriting semantic goal information—“where to go”—from large-scale visual training while retaining the robust low-level navigation behaviors learned in simulation. We propose using ordinal consistency to assess distance functions directly and demonstrate that VLD outperforms prior temporal distance approaches, such as ViNT and VIP. Experiments show that our decoupled design achieves competitive navigation performance in simulation while supporting flexible goal modalities, providing an alternative and, most importantly, scalable path toward reliable, multimodal navigation policies.
@@ 1 Introduction
Scalable training on diverse, real-world(-like), internet-scale datasets has emerged as one of the most promising approaches for learning broadly capable navigation policies. Recent advances in embodied AI have shown that policies trained on large-scale multimodal data exhibit robustness and generalization that narrow, task-specific approaches fail to achieve [NaVILA]. This highlights the importance of pursuing navigation research under conditions that better mirror the real world, rather than limiting it to controlled lab settings [roboticsChallnages, city].
Traditional imitation learning and expert demonstration pipelines offer a straightforward pathway for training navigation agents [ChauffeurNet]; however, since expert-labeled actions—typically in the form of future trajectories represented as 2D or 2.5D waypoints—are expensive to collect, they remain fundamentally constrained by the scarcity of annotated trajectories [navitrace]. Initial large-scale efforts, such as GNM [gnm], ViNT [vint], and NoMaD [nomad], have collected on the order of 200 hours of demonstrations across multiple robotic embodiments, illustrating the promise of cross-embodiment learning through curated multi-robot datasets. Nevertheless, these datasets still exhibit strong straight-walking bias (the majority of labeled actions correspond to simply moving forward), raising doubts about whether annotation-heavy pipelines can scale to broadly generalizable navigation. Moreover, such approaches largely overlook embodiment-specific signals and rely solely on image inputs, thereby precluding the use of proprioception, which is often crucial for effective navigation.
@@
Reinforcement learning (RL) enables agents to develop robustness through interaction and exploration. However, due to its sample inefficiency in real-world environments, RL training for navigation heavily depends on simulators. Collecting diverse environmental assets and simulating photorealistic images with corresponding physical interactions remain open research challenges, hindering the effective transfer of visual navigation policies from simulation to reality [10669238, roboticsChallnages, vr-robo, habitat19iccv, city].
@@
We curate  3,000 hours of diverse trajectories from synthetic environments [habitat19iccv], internet-scale walking videos [city], and embodied datasets [ego4d, scanD, GTDataset], spanning a wide variety of indoor and outdoor scenes. This design lets large-scale video learning provide the perception of where to navigate, while RL in simulation focuses on training a sensory–motor control policy for how to navigate, without relying on vision and instead using geometric signals that transfer robustly from simulation to the real world [lee2024learning]. In this way, distance prediction emerges as a central self-supervised signal for general navigation.
To broaden applicability, we extend the problem to the more general setting of vision–language goal navigation. In practice, specifying a navigation target via natural language is often more convenient than providing a reference image, especially in real-world deployments where an image may not be available. To enable this, we bootstrap a subset of our data with vision–language models [blip], generating textual goal descriptions paired with video trajectories [hirose2025omnivla]. This allows us to train a multimodal distance predictor that can operate with either image or text goals. We refer to this function as the Vision–Language Distance (VLD), which estimates the distance between the current egocentric view and a goal specified in either modality.
@@
Together, these components form the foundation of our framework for scalable, multimodal, and transferable navigation policies. Our main contributions are: (i) a scalable framework for training vision–language distance (VLD) predictors on internet-scale real-world data, outperforming prior designs for temporal distance functions such as ViNT [vint] and VIP [vip]; (ii) a novel evaluation methodology for temporal distance functions, introducing ordinal consistency as the first principled way to evaluate distance predictors in isolation rather than only through downstream policy success; and (iii) a navigation framework that decouples RL policy training in simulation from perception learning on internet-scale data, combining robust RL control with scalable semantic understanding.
@@ 2.1 Temporal Distance Functions
Temporal distance has evolved from graph-based heuristics to scalable self-supervised learning. ViNG regressed temporal steps between observation pairs with negative mining and planned over an experience graph [VING], but lacked closed-loop RL policies for robustness. Building on this idea, ViNT and NoMaD applied Transformer encoders over observation and goal embeddings, supervising both distance and short-horizon actions [vint, nomad]. While validating distance as a useful signal, these approaches primarily treated it as an auxiliary to the policy, leaving open the question of how well the distance estimator generalizes in isolation.
VIP reframed representation learning as goal-conditioned value prediction, where Euclidean distance in latent space acts as a reward [vip]. Quasimetric RL formalized multi-goal optimal values as quasimetric distances, improving sample efficiency across state- and image-based tasks [qrl, qrl2]. Yet most evaluations still emphasize downstream success, rather than direct validation of the distance function itself.
Motivated by these gaps, we adopt pairwise supervision with negative mining but design a Transformer decoder that queries distance from the goal signal. Unlike prior work, we treat the distance estimator as a first-class product, and replace mean-squared error with an inlier–outlier Gaussian mixture NLL [gaus-nill2], yielding calibrated predictions and uncertainty estimates that provide a reliability signal for downstream policies. Furthermore, we introduce a direct evaluation methodology for distance functions based on ordinal consistency, which tests whether predicted distances rank pairs in the same order as ground-truth distances—providing a principled way to validate distance estimators in isolation rather than only through downstream policy success.
@@ 2.2 Navigation Policies
Goal-conditioned RL trains policies conditioned on state, image, or language goals [GCRL]. While point-goal tasks are largely solved [point-goal], image- and language-goal navigation remain challenging due to reliance on simulators and sim-to-real gaps [vr-robo, sim2real]. To mitigate this, we train RL policies with privileged geometric distances in simulation and replace them at deployment with predictions from our learned VLD, reducing dependence on hard-to-obtain simulator photorealism for navigation training [escontrela2025gaussgym]. Related distillation and privileged-information approaches [anymalPriv, RMA, fu2022learningdeepsensorimotorpolicies] differ in that we directly substitute a raw privileged value with a learned predictor, offering a simple yet effective sim-to-real bridge.
Imitation learning from demonstrations [drivingIL, visgoalnav, gnm, vint, nomad, city] has improved generalization through curated datasets, but remains limited by annotation costs. Our approach avoids action labels entirely, scaling supervision via self-supervised distance learning on internet-scale video.
@@ 2.3 Multimodal Navigation
The emergence of vision–language–action (VLA) models [openVLA, NaVILA] has pushed toward tightly coupling perception, language understanding, and control in a single foundation model. While this integration achieves impressive generality, it comes with a high computational cost and limited scalability. In contrast, our approach explicitly decouples perception from control: we extend the temporal distance function with CLIP-based text encoding, enabling a Vision–Language Distance (VLD) predictor that handles both image and language goals. This design maintains the practicality of multimodal goals while simplifying scaling, since perception can benefit from internet-scale pretraining without requiring full-scale language models.
@@ 3.1.1 Evaluation Methodology
We evaluate this principal property with ordinal consistency, which measures whether predicted distances decrease as the agent approaches the goal and increase when it moves away. This is quantified using Kendall’s $\tau$, a rank correlation coefficient, lying in the range [1,1][-1,1], where $\tau=1$ indicates perfect agreement, $\tau=-1$ indicates complete disagreement, and $\tau=0$ corresponds to random ordering [Kendall1938-ik]. It is fully scale-invariant, capturing correlation between predicted and ground truth distances, and aligning with our focus on relative rather than absolute accuracy.
@@ 3.1.2 Model architecture
Encoders.   RGB observations are encoded with DINOv2 (small) [dino]. Text goals are processed by a CLIP text encoder (ViT-B/32) [clip], projected into the same dimensional space as DINOv2-small tokens. Both encoders remain frozen in all experiments.
@@ 3.1.3 Training Procedure
Concretely, we model the likelihood of the ground-truth temporal distance $td_{n}$ given the prediction $\hat{t}_{n}$ as a inlier–outlier Gaussian mixture [gaus-nill2]: with probability $\hat{c}_{n}$, the error is drawn from a tight “reliable” Gaussian with variance $\sigma_{R}^{2}$, and with probability $1-\hat{c}_{n}$, it comes from a broader “outlier” Gaussian with variance $\sigma_{O}^{2}$, where $\sigma_{R}\ll\sigma_{O}$. Then the training objective is the negative log-likelihood (NLL) over a batch of $N$ samples:
@@ 3.2.1 Policy Network
Inputs are fed into an LSTM core that maintains temporal memory. The LSTM output is processed by two MLP heads: a policy head that outputs navigation actions and a value head used for PPO optimization [ppo]. The noise-injected distance signal during training encourages robustness and mimics the uncertainty patterns of the VLD model.
@@ 3.2.2 Noise Function
    Relative camera rotation. The relative rotation $R_{\text{rel}}\in SO(3)$, flattened to 9 parameters [geist2024learning].
@@ 3.2.3 Reward Function
For policy training, we directly extend the Zero-Exploration Reward (ZER) formulation [zer] to explicitly encourage the agent to orient toward the goal. At each step $t$, the agent’s overlap with the goal is quantified by the projection success ratio $psr_{t}\in[0,1]$. Let $psr_{t}^{\max}$ denote the maximum overlap achievable at the current position if the agent were rotated optimally, and let $psr_{s}$ be a success threshold. The reward is then given as
@@ 4.1 Distance Function Experiments
We design experiments to demonstrate that our distance function provides a meaningful measure for navigation by producing predictions that correlate with ground-truth distances. In particular, we focus on verifying that the predicted distances have the potential to serve as a reliable navigation signal even in cases where there is no visual overlap between the agent’s current observation and the goal image.
4.1.1 Baselines
@@
The first family is based on ViNT [vint]. It provides a strong baseline due to its training on diverse real-world datasets and a comparable parameter count to our model ($\sim$31M). We evaluate both the publicly released pretrained ViNT and a fine-tuned variant on our data (ViNT-Tuned).
The second family follows VIP [vip]. The original VIP model employs a ResNet-50 backbone [resnet] and was primarily developed for manipulation tasks. For a fair comparison, we evaluate the released VIP model, a retrained navigation-oriented version (VIP-Nav), and a DINOv2-enhanced variant (VIP-DINOv2) where the ResNet encoder is replaced with a frozen DINOv2 [dino].
@@ 4.1.2 Dataset
Table 1: Ordinal consistency (Kendall’s $\tau$) on Habitat datasets.
Model HM3D (↑) Gibson (↑)
20 50 100 20 50 100
ViNT 0.40 0.35 0.28 0.53 0.41 0.39
ViNT-Tuned 0.67 0.56 0.48 0.78 0.65 0.64
VIP 0.29 0.19 0.17 0.39 0.30 0.29
VIP-Nav 0.55 0.45 0.38 0.65 0.51 0.50
VIP-DINOv2 0.42 0.45 0.40 0.62 0.52 0.51
VLD (synthetic) 0.81 0.71 0.62 0.84 0.74 0.71
VLD (real-world) 0.10 0.07 0.07 0.12 0.10 0.09
VLD (all) 0.82 0.70 0.61 0.84 0.73 0.71
4.1.3 Ordinal Consistency Evaluation
@@
Table 2: Ordinal consistency (Kendall’s $\tau$) on real-world trajectories. VLD (all) generalizes best across both settings.
Model In-the-wild (↑) Embodiment (↑)
50 100 50 100
ViNT 0.35 0.25 0.41 0.32
ViNT-Tuned 0.40 0.29 0.48 0.37
VIP 0.28 0.20 0.43 0.33
VIP-Nav 0.32 0.23 0.46 0.39
VIP-DINOv2 0.42 0.30 0.52 0.44
VLD (synthetic) 0.44 0.31 0.58 0.48
VLD (real-world) 0.23 0.18 0.16 0.14
VLD (all) 0.69 0.61 0.73 0.63
@@
Qualitative results. Across all datasets, we observe a consistent pattern: ordinal consistency is highly reliable whenever even a small amount of visual overlap exists between the current observation and the goal. This holds across both Habitat trajectories and real-world sequences (see Figures 3 and 4, with additional examples in Appendix I). The image-goal examples (3(a) and 3(b)) illustrate how VLD closely follows the ground-truth monotonic trend and reacts promptly when goal-relevant cues enter the field of view, outperforming ViNT-Tuned and VIP-Nav. Similarly, the text-goal example (Figure 4) shows that text-only predictions remain coherent—albeit noisier—and that multimodal conditioning effectively aligns with image-based performance.
In scenarios where the goal lies completely outside the field of view, confidence scores drop markedly, and the predicted distances become noisier, reflecting true perceptual uncertainty rather than model failure. This motivates analyzing whether VLD still provides a meaningful navigation signal even when no visual overlap is present. To that end, we introduce a distance accuracy metric that compares relative distances between pairs of observations without goal image visual overlap. Full details and results are provided in Appendix J.

### v2

@@ Abstract
Training end-to-end policies from image data to directly predict navigation actions for robotic systems has proven inherently difficult. Existing approaches often suffer from either the sim-to-real gap during policy transfer or a limited amount of training data with action labels. To address this problem, we introduce Vision–Language Distance (VLD) learning, a scalable framework for goal-conditioned navigation that decouples perception learning from policy learning. Instead of relying on raw sensory inputs during policy training, we first train a self-supervised distance-to-goal predictor on internet-scale video data. This predictor generalizes across both image- and text-based goals, providing a distance signal that can be minimized by a reinforcement learning (RL) policy. The RL policy can be trained entirely in simulation using privileged geometric distance signals, with injected noise to mimic the uncertainty of the trained distance predictor. At deployment, the policy consumes VLD predictions, inheriting semantic goal information—“where to go”—from large-scale visual training while retaining the robust low-level navigation behaviors learned in simulation. We propose using ordinal consistency to assess distance functions directly and demonstrate that VLD outperforms prior temporal distance approaches, such as ViNT and VIP. Experiments show that our decoupled design achieves competitive navigation performance in simulation with strong sim-to-real transfer, providing an alternative and, most importantly, scalable path toward reliable, multimodal navigation policies.
@@ 1 Introduction
Scalable training on diverse, real-world(-like), internet-scale datasets has emerged as one of the most promising approaches for learning broadly capable navigation policies. Recent advances in embodied AI have shown that policies trained on large-scale multimodal data exhibit robustness and generalization that narrow, task-specific approaches fail to achieve [4]. This highlights the importance of pursuing navigation research under conditions that better mirror the real world, rather than limiting it to controlled lab settings [51, 26].
Traditional imitation learning and expert demonstration pipelines offer a straightforward pathway for training navigation agents [2]; however, since expert-labeled actions—typically in the form of future trajectories represented as 2D or 2.5D waypoints—are expensive to collect, they remain fundamentally constrained by the scarcity of annotated trajectories [47]. Initial large-scale efforts, such as GNM [40], ViNT [41], and NoMaD [42], have collected on the order of 200 hours of demonstrations across multiple robotic embodiments, illustrating the promise of cross-embodiment learning through curated multi-robot datasets. Nevertheless, these datasets still exhibit strong straight-walking bias (the majority of labeled actions correspond to simply moving forward), raising doubts about whether annotation-heavy pipelines can scale to broadly generalizable navigation. Moreover, such approaches largely overlook embodiment-specific signals and rely solely on image inputs, thereby precluding the use of proprioception, which is often crucial for effective navigation.
@@
Reinforcement learning (RL) enables agents to develop robustness through interaction and exploration. However, due to its sample inefficiency in real-world environments, RL training for navigation heavily depends on simulators. Collecting diverse environmental assets and simulating photorealistic images with corresponding physical interactions remain open research challenges, hindering the effective transfer of visual navigation policies from simulation to reality [3, 51, 53, 36, 26].
@@
We curate  3,000 hours of diverse trajectories from synthetic environments [36], internet-scale walking videos [26], and embodied datasets [12, 17, 9], spanning a wide variety of indoor and outdoor scenes. This design lets large-scale video learning provide the perception of where to navigate, while RL in simulation focuses on training a sensory–motor control policy for how to navigate, without relying on vision and instead using geometric signals that transfer robustly from simulation to the real world [22]. In this way, distance prediction emerges as a central self-supervised signal for general navigation.
To broaden applicability, we extend the problem to the more general setting of vision–language goal navigation. In practice, specifying a navigation target via natural language is often more convenient than providing a reference image, especially in real-world deployments where an image may not be available. To enable this, we bootstrap a subset of our data with vision–language models [23], generating textual goal descriptions paired with video trajectories [14]. This allows us to train a multimodal distance predictor that can operate with either image or text goals. We refer to this function as the Vision–Language Distance (VLD), which estimates the distance between the current egocentric view and a goal specified in either modality.
@@
Together, these components form the foundation of our framework for scalable, multimodal, and transferable navigation policies. Our main contributions are: (i) a scalable framework for training vision–language distance (VLD) predictors on internet-scale real-world data, outperforming prior designs for temporal distance functions such as ViNT [41] and VIP [27]; (ii) a novel evaluation methodology for temporal distance functions, introducing ordinal consistency as the first principled way to evaluate distance predictors in isolation rather than only through downstream policy success; and (iii) a navigation framework that decouples RL policy training in simulation from perception learning on internet-scale data, combining robust and transferable RL control with scalable semantic understanding.
@@ 2.1 Temporal Distance Functions
Temporal distance has evolved from graph-based heuristics to scalable self-supervised learning. ViNG regressed temporal steps between observation pairs with negative mining and planned over an experience graph [38], but lacked closed-loop RL policies for robustness. Building on this idea, ViNT and NoMaD applied Transformer encoders over observation and goal embeddings, supervising both distance and short-horizon actions [41, 42]. While validating distance as a useful signal, these approaches primarily treated it as an auxiliary to the policy, leaving open the question of how well the distance estimator generalizes in isolation.
VIP reframed representation learning as goal-conditioned value prediction, where Euclidean distance in latent space acts as a reward [27]. Quasimetric RL formalized multi-goal optimal values as quasimetric distances, improving sample efficiency across state- and image-based tasks [45, 29]. Yet most evaluations still emphasize downstream success, rather than direct validation of the distance function itself.
Motivated by these gaps, we adopt pairwise supervision with negative mining but design a Transformer decoder that queries distance from the goal signal. Unlike prior work, we treat the distance estimator as a first-class product, and replace mean-squared error with an inlier–outlier Gaussian mixture NLL [30], yielding calibrated predictions and uncertainty estimates that provide a reliability signal for downstream policies. Furthermore, we introduce a direct evaluation methodology for distance functions based on ordinal consistency, which tests whether predicted distances rank pairs in the same order as ground-truth distances—providing a principled way to validate distance estimators in isolation rather than only through downstream policy success.
@@ 2.2 Navigation Policies
Goal-conditioned RL trains policies conditioned on state, image, or language goals [25]. While point-goal tasks are largely solved [46], image- and language-goal navigation remain challenging due to reliance on simulators and sim-to-real gaps [53, 49]. To mitigate this, we train RL policies with privileged geometric distances in simulation and replace them at deployment with predictions from our learned VLD, reducing dependence on hard-to-obtain simulator photorealism for navigation training [7]. Related distillation and privileged-information approaches [28, 21, 10] differ in that we directly substitute a raw privileged value with a learned predictor, offering a simple yet effective sim-to-real bridge.
Imitation learning from demonstrations [6, 39, 40, 41, 42, 26] has improved generalization through curated datasets, but remains limited by annotation costs. Our approach avoids action labels entirely, scaling supervision via self-supervised distance learning on internet-scale video.
@@ 2.3 Multimodal Navigation
The emergence of vision–language–action (VLA) models [19, 4] has pushed toward tightly coupling perception, language understanding, and control in a single foundation model. While this integration achieves impressive generality, it comes with a high computational cost and limited scalability. In contrast, our approach explicitly decouples perception from control: we extend the temporal distance function with CLIP-based text encoding, enabling a Vision–Language Distance (VLD) predictor that handles both image and language goals. This design maintains the practicality of multimodal goals while simplifying scaling, since perception can benefit from internet-scale pretraining without requiring full-scale language models.
@@ 3.1.1 Evaluation Methodology
We evaluate this principal property with ordinal consistency, which measures whether predicted distances decrease as the agent approaches the goal and increase when it moves away. This is quantified using Kendall’s $\tau$, a rank correlation coefficient, lying in the range [1,1][-1,1], where $\tau=1$ indicates perfect agreement, $\tau=-1$ indicates complete disagreement, and $\tau=0$ corresponds to random ordering [18]. It is fully scale-invariant, capturing correlation between predicted and ground truth distances, and aligning with our focus on relative rather than absolute accuracy.
@@ 3.1.2 Model architecture
Encoders.   RGB observations are encoded with DINOv2 (small) [33]. Text goals are processed by a CLIP text encoder (ViT-B/32) [34], projected into the same dimensional space as DINOv2-small tokens. Both encoders remain frozen in all experiments.
@@ 3.1.3 Training Procedure
Concretely, we model the likelihood of the ground-truth temporal distance $td_{n}$ given the prediction $\hat{t}_{n}$ as a inlier–outlier Gaussian mixture [30]: with probability $\hat{c}_{n}$, the error is drawn from a tight “reliable” Gaussian with variance $\sigma_{R}^{2}$, and with probability $1-\hat{c}_{n}$, it comes from a broader “outlier” Gaussian with variance $\sigma_{O}^{2}$, where $\sigma_{R}\ll\sigma_{O}$. Then the training objective is the negative log-likelihood (NLL) over a batch of $N$ samples:
@@ 3.2.1 Policy Network
Inputs are fed into an LSTM core that maintains temporal memory. The LSTM output is processed by two MLP heads: a policy head that outputs navigation actions and a value head used for PPO optimization [37]. The noise-injected distance signal during training encourages robustness and mimics the uncertainty patterns of the VLD model.
@@ 3.2.2 Noise Function
    Relative camera rotation. The relative rotation $R_{\text{rel}}\in SO(3)$, flattened to 9 parameters [11].
@@ 3.2.3 Reward Function
For policy training, we directly extend the Zero-Exploration Reward (ZER) formulation [1] to explicitly encourage the agent to orient toward the goal. At each step $t$, the agent’s overlap with the goal is quantified by the projection success ratio $psr_{t}\in[0,1]$. Let $psr_{t}^{\max}$ denote the maximum overlap achievable at the current position if the agent were rotated optimally, and let $psr_{s}$ be a success threshold. The reward is then given as
@@ 4.1 Distance Function Experiments
Table 1: Ordinal consistency (Kendall’s $\tau$) across Habitat and real-world datasets.
Model Habitat (↑) Real-world (↑)
HM3D Gibson In-the-wild Embodiment
20 50 100 20 50 100 50 100 50 100
ViNT 0.40 0.35 0.28 0.53 0.41 0.39 0.35 0.25 0.41 0.32
ViNT-Tuned 0.67 0.56 0.48 0.78 0.65 0.64 0.40 0.29 0.48 0.37
VIP 0.29 0.19 0.17 0.39 0.30 0.29 0.28 0.20 0.43 0.33
VIP-Nav 0.55 0.45 0.38 0.65 0.51 0.50 0.32 0.23 0.46 0.39
VIP-DINOv2 0.42 0.45 0.40 0.62 0.52 0.51 0.42 0.30 0.52 0.44
VLD (synthetic) 0.81 0.71 0.62 0.84 0.74 0.71 0.44 0.31 0.58 0.48
VLD (real-world) 0.10 0.07 0.07 0.12 0.10 0.09 0.23 0.18 0.16 0.14
VLD (all) 0.82 0.70 0.61 0.84 0.73 0.71 0.69 0.61 0.73 0.63
We design experiments to demonstrate that our distance function provides a meaningful measure for navigation by producing predictions that correlate with ground-truth distances. In particular, we focus on verifying that the predicted distances can serve as a reliable navigation signal even when there is no visual overlap between the agent’s current observation and the goal image.
4.1.1 Baselines.
@@
The first family is based on ViNT [41]. It provides a strong baseline due to its training on diverse real-world datasets and a comparable parameter count to our model ($\sim$31M). We evaluate both the publicly released pretrained ViNT and a fine-tuned variant on our data (ViNT-Tuned).
The second family follows VIP [27]. The original VIP model employs a ResNet-50 backbone [13] and was primarily developed for manipulation tasks. For a fair comparison, we evaluate the released VIP model, a retrained navigation-oriented version (VIP-Nav), and a DINOv2-enhanced variant (VIP-DINOv2) where the ResNet encoder is replaced with a frozen DINOv2 [33].
@@
4.1.3 Ordinal Consistency Evaluation.
@@
Qualitative results. Across all datasets, we observe a consistent pattern: ordinal consistency is highly reliable whenever even a small amount of visual overlap exists between the current observation and the goal. This holds across both Habitat trajectories and real-world sequences (see Figure 3 and 4, with additional examples in Appendix I). The image-goal examples (3(a) and 3(b)) illustrate how VLD closely follows the ground-truth monotonic trend and reacts promptly when goal-relevant cues enter the field of view, outperforming ViNT-Tuned and VIP-Nav. Similarly, the text-goal example (Figure 4) shows that text-only predictions remain coherent—albeit noisier—and that multimodal conditioning effectively aligns with image-based performance.
In scenarios where the goal lies completely outside the field of view, confidence scores drop markedly and the predicted distances become noisier, reflecting true perceptual uncertainty rather than model failure. This motivates analyzing whether VLD still provides a meaningful navigation signal even when no visual overlap is present. To that end, we introduce a distance accuracy metric that compares relative distances between pairs of observations without goal image visual overlap. Full details and results are provided in Appendix J.